A man built a walk of uniform width around a rectangular pool. If the area of the walk is 145 square feet and the dimensions of the pool are 17 feet by 7 feet, how wide is the walk?
Answer by richwmiller(17219)

(x+17)*(x+7)-17*7=145
(x+7)*(x+17)-119 = 145
x^2+24x-145 = 0
x=-29 disregard negative answers
x = 5